OpenStudy (anonymous):

-4x+3=6x-7...x=-2??? Can someone c if this the right answer

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

You ned to get rid of the -4x first by adding 4x to both sides.

OpenStudy (anonymous):

3 = 10x - 7

OpenStudy (anonymous):

add 7 to both sides

OpenStudy (anonymous):

10 = 10x

OpenStudy (anonymous):

divide both sides by 10

OpenStudy (anonymous):

x = 1

OpenStudy (anonymous):

so i get rid of the - sign in front of the -4x and just add it to the 6x

OpenStudy (anonymous):

Yes you get rid of the -4x by +4x to both sides. I get rid of the -4x because it is smaller than 6x...

OpenStudy (anonymous):

ok cuz i used-4 + 6x and got -2x im lost

OpenStudy (anonymous):

To get rid of let's say +5 you would -5 because that would give you 0 To get rid of -3 you + 3 because that would give you 0 To get rid of +7x you -7x because that would give you 0 To get rid of -4x you +4x because that would give you 0 Does that help?

OpenStudy (anonymous):

actually since u explained it like that yes it does......sry if im being high maintenance hate not understanding my teacher sucks at his job

OpenStudy (anonymous):

x/-7-9 =-3......the / means divide so i need to multiply?

OpenStudy (anonymous):

OK.. You only have an x on one side so you need to get the x by itself. You always add or subtract first if needed. So you need to get rid of the -9 first

OpenStudy (anonymous):

x/-7=-12?

OpenStudy (anonymous):

To get rid of -9 you need to +9 -3 + 9 = 6 so x/-7 = 6 Do you understand?

OpenStudy (anonymous):

not really.....

OpenStudy (anonymous):

When it says to solve.. The first thing you should do is go to the side with the "x".. If both sides of the = sign have an "x" then you get rid of the smallest "x" by either adding or subtracting the "x". you do the opposite, if it is a positive "x" then you subtract. if it is a negative "x" then you add.. In this case, you only have x on one side. Do you know how to do this problem. x + 3 = 10

OpenStudy (anonymous):

x=7

OpenStudy (anonymous):

Without actually just figuring that out... do you know what to do algebraically?

OpenStudy (anonymous):

yes subtract 3 from both sides

OpenStudy (anonymous):

Good!!! Look at the problem we are doing and pretend the -7 is not there... we would have x - 9 = -3 How would you solve this?

OpenStudy (anonymous):

add 9 to both sides

OpenStudy (anonymous):

Good.. That is what you do first on this problem. NOw.. How would you solve 2x = 10 (algebraically)

OpenStudy (anonymous):

divide 10 by 2 since 2x is multiplying u need to divide

OpenStudy (anonymous):

Great.. So let's do our problem... x/-7 - 9 = -3 You add 9 to both sides x/-7 = 6 then to get rid of "divide by -7" you multiply by -7 .. put a (-7) to the left of the x and to the right of the 6 (-7) x = 6 (-7) ----- -7

OpenStudy (anonymous):

the -7's cancel and you get x = -42

OpenStudy (anonymous):

making sense so if x/-6=-4 can i use the the -4 or the -6? sry if stupid ?

OpenStudy (anonymous):

You always go to the side with the x... don't even look at the other side. So all you see is x/-6 What would you do then?

OpenStudy (anonymous):

add 6 so x=2

OpenStudy (anonymous):

Read this to yourself x/-6 Did you say x divided by -6?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So to get rid of divided by -6 you should multiply by -6. (-6)x -4(-6) ----- = -6 The -6's cancel and you get x = 24

OpenStudy (anonymous):

wow im retarded got frazzled since there was only 2 numbers.......

OpenStudy (anonymous):

Do you have another problem?

OpenStudy (anonymous):

im sure i will got like 20 left for homework will u b on for a while

OpenStudy (anonymous):

Sure if you want me to help you.

OpenStudy (anonymous):

thank u so much for taking time and being paitent

OpenStudy (anonymous):

real quick lol 3x+5=4x+16......x=3 or -11=x

OpenStudy (anonymous):

x = -11 3x + 5 = 4x + 16 get rid of your smallest x (3x) by subtracting 3x from both sides 5 = x + 16 subtract 16 from both sides -11 = x x = -11

OpenStudy (anonymous):

thats what i got my friend and i were in a debate

OpenStudy (anonymous):

thx again

OpenStudy (anonymous):

do u know how to use a algebraic calculator

OpenStudy (anonymous):

Probably.. what do you need to know

OpenStudy (anonymous):

trying to check my work not sure how to enter the equation my friend gave me a ti 30xa

OpenStudy (anonymous):

You mean of the problem we just did? Just type 3(-11) + 5 = You should see -28 Now type 4(-11)+16 = You should see -28 That means you were correct

OpenStudy (anonymous):

dont i need to put x in the equation

OpenStudy (anonymous):

No ... you put in what you got for x... in the last problem it was -11

OpenStudy (anonymous):

ok -2(3x+1)=3x-20......can we do together?

OpenStudy (anonymous):

Yes. When you have parentheses, you need to get rid of them first... you do this by distributive method.. can you do this?

OpenStudy (anonymous):

yes -6x + -2=3x-20....first step

OpenStudy (anonymous):

Good... It's best not to have a + next to a - Just write -2 * 3x = -6x and -2* 1 = -2 So we write -6x - 2 = 3x - 20 NOW you have x's on both sides - so get rid of the smallest x.

OpenStudy (anonymous):

so y do we change from adding to subtracting

OpenStudy (anonymous):

WOW.. hard to explain... Were you ever taught that when you were given a "- 5" you could change that into a "+ a negative"

OpenStudy (anonymous):

Plus a negative 2 (+ -2) is the same thing as subtracting 2 (-2)

OpenStudy (anonymous):

You just never want two signs next to each other... So going back to the problem -2(3x + 1) you should think of this. -2(3x) and -2(+1) so you get -6x -2 **** THE SIGN IN FRONT OF A NUMBER ALWAYS GOES WITH THAT NUMBER... it's like it belongs to that number.

OpenStudy (anonymous):

We can do it your way... here we go -6x + -2 = 3x - 20 How do we get x's on one side (get rid of the smallest "x")

OpenStudy (anonymous):

ur way makes alot more sense here we go

OpenStudy (anonymous):

just to make sure if i use the smaller x it would be 9x?

OpenStudy (anonymous):

very good -2 = 9x - 20 NOW that you have x on one side ONLY LOOK AT THE side with the x... What do you get rid of next.

OpenStudy (anonymous):

or - 3x

OpenStudy (anonymous):

To get rid of the -6x you need to +6x to both sides.

OpenStudy (anonymous):

add 20 or 2 to both sides get -18.....-18/9x= x=-2?

OpenStudy (anonymous):

Very important you understand what I am saying. When there is only one x in your problem. You go to the side of the equal sign that has the x. YOUR EYES SHOULD NOT EVEN SEE THE OTHER SIDE OF THE = sign. In this case.. take your left hand and cover up the -2. Your eyes should ONLY see 9x - 20 and in this case you do + 20 to both sides.

OpenStudy (anonymous):

okok i knew that trying to rush so im not holding u up......x=2

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

thx again whats ur name im keith

OpenStudy (anonymous):

Andrea

OpenStudy (anonymous):

andrea i cant thk u enough

OpenStudy (anonymous):

Do you have another problem? By the way, if you ever need help after tonight.. I would be glad to help you.. you can email me at amcbeth328@gmail.com

OpenStudy (anonymous):

4(3x-5)=-4x+27 that equals 12x-20=-4x+27, so would it b 12x - -4x or 12x-4x?

OpenStudy (anonymous):

12x - 20 = -4x + 27 You have x's on both sides of the = sign.. get rid of the smallest x by doing the opposite.

OpenStudy (anonymous):

just wanted to know if next step = 8x or 16x

OpenStudy (anonymous):

-4x is smaller than 12x so you get rid of the -4x by +4x to both sides 16x - 20 = 27

OpenStudy (anonymous):

then add 20 to both

OpenStudy (anonymous):

Yes and you won't get a nice answer

OpenStudy (anonymous):

16x = 47 divide by 16 x = 2.9 or 47/16 which = 2 15/16

OpenStudy (anonymous):

Any more?

OpenStudy (anonymous):

7(4-x)+3=17......x=-2?

OpenStudy (anonymous):

28 - 7x + 3 = 17 -7x + 31 = 17 -7x = -14 x = 2 Almost.. how do we differ

OpenStudy (anonymous):

i meant x=2 my brain is fried forgot when u divide 2 negatives u get positive

OpenStudy (anonymous):

Any more? I am getting tired and need to get to bed. I have to be to school early tomorrow. But if we are close, I will hang in there.

OpenStudy (anonymous):

no im good thk u very much

OpenStudy (anonymous):

OK Keith.. keep working hard... it will pay off for you in the long run... I could see a lot of improvement with your math tonight. Have a good week at school.

OpenStudy (anonymous):

u too